OpenStudy (anonymous):

What is the solution set for |2x – 6| – 7 = 7?

hero (hero):

|2x – 6| – 7 = 7 |2x – 6| = 14 2x - 6 = 14 2x-6 = -14 2x = 20 2x = -8 x = 10 x = -4

OpenStudy (anonymous):

damn your good